# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ F Implementation Details on the QH9 Benchmark Series
The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series. When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^4] https://github.com/divelab/AIRS provided by Yu et al. (2023b), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules. We opt to apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature. We set the number of channels for $\textbf{u}^{(k)}$ $(1\leq k\leq 2)$ as $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases. The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Regarding the error metric in the loss function Eq. 7, for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$. Other hyper-parameters and configurations are the same as QHNet: the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is selected as AdamW, and a learning rate scheduler is implemented. To be more specific, the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps. Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
@@ G Visualization of Block-level MAE Statistics
As shown in Fig. 2, each Hamiltonian matrix consists of numerous basic block, with each basic block representing the direct product of two degrees. Here, we follow Yin et al. (2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$. The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 5 and 6. Fig. 5 presents the results for monolayer structures, while Fig. 6 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, DeepH-E3, across the vast majority of blocks of the Hamiltonian matrices, particularly on blocks where DeepH-E3 struggles with lower accuracy.
@@
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$).
H Ablation Study
@@
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression. In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
@@
Table 6: Ablation study MAE results (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Table 7: Ablation study MAE results (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Experimental results of the four setups are listed in Table 6 and 7. Table 6 presents the results for monolayer structures, while Table 7 focuses on bilayer structures. From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the non-linear expressiveness of SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
I Discussion on Time Costs
When the atomic system is small and any two atoms are within the cutoff radius of each other such that the Hamiltonian is defined, the number of Hamiltonian matrices is approximately proportional to $N^{2}$, where $N$ is the number of atoms in the system. Yet, as the atomic system grows, the system exhibits a dual nature of locality and non-locality. Locality refers to the fact that, when $N$ becomes large, the number of atoms within the cutoff radius of each atom remains roughly constant, and thus the growth in the number of Hamiltonian matrices asymptotically scales as $O(N)$. On the other hand, the non-locality arises from the fact that, although Hamiltonian matrices are defined only for pairs of atoms within each other’s cutoff radius, atoms outside the cutoff radius may still exert long-range effects on the Hamiltonian matrices within the cutoff radius
The baseline models we select, whether DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b), are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms, which cleverly balance the locality and non-locality of the system. As a result, the computational complexity asymptotically scales as $O(N)$ as $N$ grows to large. This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT methods, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems. Our method improves the Hamiltonian prediction accuracy of DeepH-E3 and QHNet while maintaining this advantage of asymptotic complexity. The neural module $g_{nonlin}(\cdot)$ we introduce efficiently updates each edge feature $\textbf{f}^{(k)}$ $(1\leq k\leq K)$ from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature $\textbf{v}^{(k)}$, ensuring that the computational complexity remains asymptotically $O(N)$, with good potential to handle large atomic systems at acceptable computational costs.
We test the inference time costs of the DeepH-E3 and DeepH-E3+TraceGrad models on an Nvidia RTX A6000 GPU, using the testing sets from two databases, namely $MG$ and $MM$, as examples; and test the inference time costs of the QHNet and QHNet+TraceGrad models on $QS$ and $QD$, on the same device. The batch size is set to 1 for all tests on inference time. Each sample is inferred 50 times, and the average time per inference is recorded. When using DeepH-E3, the average inference time per sample (s/sample) for $MG$ and $MM$ are as follows: $0.247$ and $0.256$, respectively; with DeepH-E3+TraceGrad, the average inference times are $0.264$ and $0.274$, correspondingly. When using QHNet, the average inference time per sample (s/sample) for $QS$ and $QD$ are as follows: $0.233$ and $0.174$, respectively; with QHNet+TraceGrad, the average inference times are $0.248$ and $0.187$, correspondingly. As we can see, the increases in computational costs due to the TraceGrad module are slight relative to that of the baseline models. Considering the significant accuracy improvements brought by the TraceGrad method as reported in Section 6.2, we believe that this minor increase in computational cost is acceptable and worthwhile in practical applications.

### v5

@@ F.2 Implementation Details on the QH9 Benchmark Series
The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series. When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^3] https://github.com/divelab/AIRS provided by Yu et al. (2023b), keeping the architecture and network configurations consistent with their setup. In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules. We opt to apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature. We set the number of channels for $\textbf{u}^{(k)}$ $(1\leq k\leq 2)$ as $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases. The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for an atomic pairwise Hamiltonian matrix. Regarding the error metric in the loss function Eq. (3), for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$. Other hyper-parameters and configurations are the same as QHNet: the maximum training steps are set as $300,000$ for $QS$ and $260,000$ for $QD$, the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is set as AdamW, and a learning rate scheduler is implemented: the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps. Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
@@ G Visualization of Block-level MAE Statistics
As shown in Fig. 2, each Hamiltonian matrix consists of numerous basic blocks, with each basic block representing the direct product of two degrees. Here, we follow (Yin et al., 2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$. The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 5 and 6. Fig. 5 presents the results for monolayer structures, while Fig. 6 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, DeepH-E3, across the vast majority of blocks of the Hamiltonian matrices, particularly on blocks where DeepH-E3 struggles with lower accuracy.
@@
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases.
H Ablation Study and Comparison between the Proposed Gradient-based Mechanism with the Gated Activation Mechanism
@@
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{H}\cdot\mathbf{H}^{\dagger})=\text{tr}(\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}\cdot(\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by the SO(3)-equivariant encoders of DeepH-E3 for Hamiltonian regression. In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
@@
Table 5: MAE results (meV) of multiple experimental settings on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Table 6: MAE results (meV) of multiple experimental settings on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Furthermore, to compare our proposed gradient-based mechanism with the gated activation mechanism (Weiler et al., 2018) which is widely used in SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023), the following experimental setups are also included:
• DeepH-E3+Gate: This variant modifies the experimental setup from DeepH-E3+Grad by replacing the gradient mechanism, which constructs equivariant features as $\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$, with the gated activation mechanism, constructing equivariant features as $\mathbf{v}=z\cdot\mathbf{f}$. All other aspects remain the same.
• DeepH-E3+TraceGate: This variant modifies the experimental setup from DeepH-E3+TraceGrad by replacing the gradient mechanism ($\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$) with the gated activation mechanism ($\mathbf{v}=z\cdot\mathbf{f}$). All other aspects remain the same.
Experimental results for all of the six setups are listed in Table 5 and 6. Table 5 presents the results for monolayer structures, while Table 6 focuses on bilayer structures. We have the following observations:
@@
First, by comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation level, can contribute to the performance individually. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the non-linear expressiveness of SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
Second, it is evident that both DeepH-E3+Gate and DeepH-E3+TraceGate show improvement over the baseline DeepH-E3, as these configurations introduce an additional neural network module $s_{nonlin}(\cdot)$ in learning the feature $z$, which increases model capacity. Moreover, DeepH-E3+TraceGate also incorporates our proposed trace supervision signal, which guides the learning of SO(3)-invariant features. However, their performance falls short of DeepH-E3+Grad and DeepH-E3+TraceGrad, respectively, indicating that the gated activation mechanism may not fully capture the system’s non-linearity patterns, whereas the proposed gradient-based mechanism demonstrates stronger generalization performance on Hamiltonian prediction and may be a better choice in terms of expressive capability.
I Theoretical Analysis on Computational Complexity
The total number of non-zero Hamiltonian matrix elements to be calculated is proportional to the number of local atomic pairs in the system, with a complexity of $\mathcal{O}(N\overline{E})$, where $N$ is the total number of atoms and $\overline{E}$ is the average number of neighboring atoms within the cutoff radius per atom. Since the atomic orbital basis set has a finite range, the non-zero matrix elements vanish beyond a certain distance for an given atom. In small systems, where all atoms lie within each other’s cutoff radius, $\overline{E}$ scales with $N$, resulting in a total number of non-zero elements proportional to $N^{2}$. However, in sufficiently large systems, the finite range of the atomic orbitals ensures that $\overline{E}$ remains constant, independent of $N$. As a result, for large atomic systems, the total number of non-zero elements simplifies to $\mathcal{O}(N)$.
The baseline models we select, whether DeepH-E3 or QHNet, are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms. These models cleverly balance the locality of Hamiltonian definitions with the long-range interactions present in the system. As a result, the computational complexity asymptotically scales as $\mathcal{O}(N)$ as $N$ increases, which is consistent with the growth of the scales of non-zero Hamiltonian matrix elements. The proposed TraceGrad method directly updates each SO(3)-equivariant feature of the baseline models, and the computational amount is proportinal to the number of features of the baseline models. Therefore, combining TraceGrad, the computational complexity also scales as $\mathcal{O}(N)$.
Traditional DFT methods require $T$ iterations of diagonalizing $N\times N$ matrices, each with a time complexity of $\mathcal{O}(N^{3})$, because all occupied states are needed to compute the charge density. As $N$ increases, this cubic complexity leads to significant computational overhead, making it challenging to simulate large atomic systems within a reasonable time frame. In contrast, our deep learning framework enables the efficient and accurate prediction of the Hamiltonian for large atomic systems with a linear time complexity of $\mathcal{O}(N)$, eliminating the need for self-consistent iterations. Moreover, since most physical properties, such as transport, optical, and topological properties, depend only on the energy bands near the Fermi level, it is unnecessary to solve for the eigenfunctions of all occupied states once the Hamiltonian is known. Since the Hamiltonian matrix is sparse and only a limited number of bands near the Fermi level are needed, these eigenstates can be efficiently computed using methods like the shift-invert approach available in the ARPACK package (Lehoucq et al., 1998), with a computational complexity of $\mathcal{O}(N)$.
J A Joint Discussion on GPU Time Costs and Performance Gains
We here provide a joint comparison of the GPU time cost and corresponding accuracy of different models across four databases as representative: Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), QH9-stable ($QS$), and QH9-dynamic ($QD$). This comparison includes the average inference time per sample for each model, with the test batch size set as 1 and hardware environment set as Nvidia RTX A6000 GPU in single-task mode without computational sharing with other processes.
For $MG$ and $MM$, we compare among DeepH-E3, DeepH-E3+TraceGrad, and DeepH-E3×2, where DeepH-E3×2 refers to a model obtained by doubling the number of encoding blocks in DeepH-E3 and training it from scratch until convergence. For $QS$ and $QD$, we compare among QHNet, QHNet+TraceGrad, and QHNet×2, where QHNet×2 refers to a model obtained by doubling the number of encoding blocks in QHNet and training it from scratch until convergence. The experimental results are documented in the Table 7.
Table 7: Average inference time (denoted as Time, in seconds) and MAE performance ($MAE^{H}_{all}$, in meV) for testing samples from the Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), QH9-stable ($QS$), and QH9-dynamic ($QD$) databases. $\downarrow$ indicates that lower values of the metrics correspond to better accuracy. All models are tested individually on a single Nvidia RTX A6000 in single-task mode.
Methods $MG$ $MM$
$Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ ) $Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ )
DeepH-E3 (Baseline) 0.247 0.251 0.256 0.406
DeepH-E3×2 0.483 0.244 0.510 0.387
DeepH-E3+TraceGrad 0.264 0.175 0.274 0.285
Methods $QS$ $QD$
$Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ ) $Time$ ()($\downarrow$)( ↓ ) $MAE^{H}_{all}$ ()($\downarrow$)( ↓ )
QHNet (Baseline) 0.233 1.962 0.174 4.733
QHNet×2 0.497 1.845 0.385 4.532
QHNet+TraceGrad 0.248 1.191 0.187 2.819
From this Table, we find that adding the TraceGrad module results in only a slight increase in inference time compared to the baseline models. Given the substantial accuracy improvements introduced by the TraceGrad method, we consider this minor increase in computational time acceptable for practical applications. In contrast, simply increasing the depth of DeepH-E3 or QHNet results in a significant rise in inference time but yields only limited accuracy improvements. In contrast, DeepH-E3+TraceGrad demonstrates significantly better accuracy performance than DeepH-E3×2, and similarly, QHNet+TraceGrad achieves notably higher accuracy than QHNet×2. Furthermore, the inference time of DeepH-E3+TraceGrad and QHNet+TraceGrad are both lower than their respective DeepH-E3×2 and QHNet×2 counterparts. These results underscore the superiority of the TraceGrad method in enhancing expressive capability and improving accuracy performance, while maintaining time efficiency.
K Acceleration Performance for the Convergence of Traditional DFT Algorithms
Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations. In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms. We evaluate the acceleration performance brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018). Specifically, we adopt the two groups of metrics on acceleration performance:
The first group of metrics are defined by Yu et al. (2023a), as follows:
  • Achieved ratio. This metric calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e.
  • Error-level ratio. This metric measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
Table 8: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation. Both models are evaluated on a set of $50$ molecules chosen by (Yu et al., 2023a), with the mean and standard deviation of the metrics across these samples reported. $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
Methods Training DFT Metric Ratio
databases initialization
QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
Error-level ratio $\uparrow$ $0.620\pm 0.037$
minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
Error-level ratio $\uparrow$ $0.406\pm 0.021$
QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
Error-level ratio $\uparrow$ $0.622\pm 0.048$
minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
Error-level ratio $\uparrow$ $0.406\pm 0.066$
Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$
Table 9: Average wall time costs per sample (/s) for three experimental settings: $t1$ represents the wall time required for DFT calculation with a random initial guess like 1e of minao; $t2$ denotes the wall time for inference of deep learning methods (i.e., QHNet or QHNet+TraceGrad); $t3$ is the total time of the process including deep learning inference and the DFT calculation that uses the outputs of deep learning methods as initialization. Both models are evaluated on a set of $50$ molecules chosen by (Yu et al., 2023a), with the mean and standard deviation of the metrics across these samples reported. $\downarrow$ indicates that lower values correspond to better performance. In order to ensure fairness in comparison, all experimental settings including DFT calculation and deep learning inference are measured on a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU in single-task mode.
Methods Training DFT Metric Time
databases initialization
QHNet (Baseline) QS 1e t1 $120.896\pm 9.134$
t2 $\downarrow$ $1.724\pm 0.025$
t3 $\downarrow$ $48.604\pm 7.741$
minao t1 $63.193\pm 5.335$
t2 $\downarrow$ $1.724\pm 0.025$
t3 $\downarrow$ $48.604\pm 7.741$
QD 1e t1 $87.161\pm 12.075$
t2 $\downarrow$ $1.280\pm 0.019$
t3 $\downarrow$ $44.146\pm 9.342$
minao t1 $51.396\pm 5.870$
t2 $\downarrow$ $1.280\pm 0.019$
t3 $\downarrow$ $44.146\pm 9.342$
QHNet+TraceGrad QS 1e t1 $\downarrow$ $120.896\pm 9.134$
t2 $\downarrow$ $1.852\pm 0.020$
t3 $\downarrow$ $\bm{41.941\pm 6.783}$
minao t1 $63.193\pm 5.335$
t2 $\downarrow$ $1.852\pm 0.020$
t3 $\downarrow$ $\bm{41.941\pm 6.783}$
QD 1e t1 $87.161\pm 12.075$
t2 $\downarrow$ $1.361\pm 0.010$
t3 $\downarrow$ $\bm{39.712\pm 9.076}$
minao t1 $51.396\pm 5.870$
t2 $\downarrow$ $1.361\pm 0.010$
t3 $\downarrow$ $\bm{39.712\pm 9.076}$
Experimental results on these metrics are recorded in Table 8, where the results for the compared method QHNet are taken from Yu et al. (2023a), while the results of QHNet+TraceGrad, come from our experiments. In our experiments, the DFT calculation settings follow those in Section 4 of Yu et al. (2023a) (the DFT parameters and the $50$ testing samples) , except for the CPU environment, where we use a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU in single-task mode. It is worth noting that this does not affect the fairness of the comparison, as the achieved ratio and error-level ratio measure the ratio of iteration counts rather than runtime, and differences in CPU computation times are negligible in these metrics. From Table 8, we could observe that the proposed TraceGrad method brings significant improvements to the baseline model QHNet on the acceleration ratio of DFT calculation, notably reducing the achieved ratio and enhancing the error-level ratio.
The second group of metrics measures the wall time savings that deep learning methods contribute to DFT calculations, specifically quantifying the incremental time savings brought by our proposed TraceGrad method in accelerating DFT calculations. For a fair comparison, we report the average wall time costs of testing samples (/s) across three metrics:
  • $t1$: The wall time required for a DFT calculation initialized with a random guess initialization, such as 1e or minao.
  • $t2$: The wall time for inference using deep learning methods (i.e., QHNet or QHNet+TraceGrad).
  • $t3$: The total wall time for the combined process, including both deep learning inference and the subsequent DFT calculation initialized with the deep model’s outputs. $t3$ here provides a more comprehensive evaluation of the actual time savings achieved when incorporating deep learning methods.
Experimental results on these metrics are recorded in Table 9. Here all time-related measurements are conducted on a single thread of an Intel(R) Xeon(R) Gold 6330 @ 2.00GHz CPU, including the experiments for QHNet, which are reproduced under the same conditions to ensure fairness. Unlike the time measurements in Appendix J, which are performed on GPUs, all deep learning models here are evaluated on the CPU thread to maintain consistency in the comparison with DFT software. From the experimental results, we observe three key findings:
  • Comparing $t2$ and $t1$, deep models are significantly faster than DFT calculations, achieving speeds tens of times greater for the testing samples. It is worth noting that, given that the testing samples here are all small molecular systems, deep models have already demonstrated a significant time efficiency advantage compared to DFT. Based on the computational complexity analysis of deep learning methods compared to DFT methods in Appendix I, it is reasonable to infer that for larger atomic systems, the disparity between $t2$ and $t1$ will expand rapidly.
  • Comparing $t2$ values between QHNet and QHNet+TraceGrad, the additional runtime introduced by TraceGrad is relatively minor on the CPU, consistent with the GPU-based results reported in Appendix J.
  • Comparing $t3$ and $t1$, the total runtime of using a deep model to predict initial values and then running DFT calculations is significantly lower than performing DFT calculations from a random guess initialization. Particularly, comparing row 4 and row 16, row 7 and row 19, row 10 and row 22, as well as row 13 and row 25 in Table 9, it can be observed that combing TraceGrad further reduces $t_{3}$, demonstrating that the time saved by TraceGrad in DFT calculations far exceeds the minimal additional time required for its inference.